NvidiaIntely (nvidiaintely):

The half-life of iodine-131 is 8.10 days. How long will it take for three-fourths of a sample of iodine-131 to decay? 4.05 days 8.10 days 16.2 days 32.4 days

OpenStudy (kevin):

3/4 is decay, that's mean the remaining of iodine is : 1 - 3/4 = 1/4 Nt == 1/4 No Nt/No = 1/4 Nt/No = 1/2^(t/ delta t) 1/4 = (1/2)^(t/8.1) t = 16.2

NvidiaIntely (nvidiaintely):

Thanks

OpenStudy (kevin):

yw